OpenStudy (anonymous):

WRITE THE RATE AS A FRACTION IN LOWEST TERMS. 28 LETTERS FOR 6 MINUTES.. THE RATE IS .... (TYPE A SIMPLIFIED FRACTION) TYPE THE UNITS RIGHT OF THE NUMBER

OpenStudy (anonymous):

28 / 6

OpenStudy (anonymous):

i think

Mehek (mehek14):

you can simplify that

Mehek (mehek14):

look at how many times 6 goes into 28

OpenStudy (anonymous):

4 TIMES

OpenStudy (anonymous):

PLUS 4 REMAINS

Mehek (mehek14):

ok so \(4\dfrac{4}{6}\)

Mehek (mehek14):

you can simplify 4/6

Mehek (mehek14):

divide both of them by 2

OpenStudy (anonymous):

2/3

Mehek (mehek14):

\(4\dfrac{2}{3}\) is your answer

OpenStudy (anonymous):

BUT THERES NO WHOLE NUMBERS THE GIVEN IS 2 FRACTION BOXES

Mehek (mehek14):

ok so back to \(\dfrac{28}{6}\)

Mehek (mehek14):

divide both by 2 you get \(\dfrac{14}{3}\)

Mehek (mehek14):

that's the simplest form in improper fraction form

OpenStudy (anonymous):

IM LOST